1 find net torque and net force on the pulley. R:.05 FIVE STAR Q malkg m = 95 kg FIVE STAR
Answer #1

M M2 Now T = M, G - 1x 9.8 - 9.8N Tq = M, g = 0.9589.8W = 9.310 so net torce F = T, -Tz = 9.8-9.31 = 0.49N Ned torque T = 9 -